# Evaluating Semantic and Syntactic Understanding in Large Language Models for Payroll Systems

### v1

@@ 1 Introduction
Large language models (LLMs) have moved quickly from research demos to everyday tools. They follow instructions, summarize long documents, and support search and analysis at scale. With each generation, they show a stronger grasp of meaning in context, which makes them attractive for business workflows that depend on careful reading and consistent decisions. This momentum has encouraged organizations to pilot LLMs across back-office operations, customer support, and analytics. The same momentum also raises basic questions about reliability. High-stakes applications requires clear standards for transparency, accountability, and risk management, and there is ongoing evidence that fluent outputs can still be incorrect NIST2023RMF ; EUAIAct2024 ; Ji2023HallucinationSurvey ; Zeng2024AHE .
@@ 2 Literature Review
Early evaluations showed that strong LLMs still make arithmetic and logic mistakes on word problems. The GSM8K and MATH benchmarks made these issues visible and reproducible Cobbe2021GSM8K ; Hendrycks2021MATH . Follow-up studies improved results by asking models to show their steps. Chain-of-Thought (CoT) prompting and Self-Consistency reduce single-path failures and often raise accuracy Wei2022CoT ; Wang2022SelfConsistency . Another line of work reduces mistakes by letting models write and run code, or by calling tools. Program-Aided LM (PAL) and Program-of-Thoughts separate natural language reasoning from exact computation. This improves faithfulness on math and finance-style questions Gao2022PAL ; Chen2022PoT . Domain-adapted models such as Minerva also help on STEM problems, but they still benefit from careful instructions and structure Lewkowycz2022Minerva .
@@
Reasoning strategies continue to evolve. Least-to-Most prompting breaks a problem into smaller parts, which can help generalization to harder cases Zhou2022LeastToMost . Prompt engineering practices have become more systematic, including patterns that set roles, constraints, and output formats. Toolformer further shows that models can learn when to call external tools such as calculators and search Schick2023Toolformer . Recent work also probes arithmetic reliability directly and designs tests that expose subtle numerical errors, which is important when results must be exact Sun2025ArithmeticErrors ; talburt2026casecountmetriccomparative.
@@
Beyond raw reasoning, many practical tasks require models to align text with tables, spreadsheets, or databases. This alignment is critical in payroll, where each field has a specific meaning and must be used in the right order. In table question answering, TaPas learns table representations that support QA without explicit logical forms, and TAPEX pretrains a neural executor for table operations Herzig2020TaPas ; Liu2021TAPEX . For database-backed settings, large text-to-SQL datasets such as WikiSQL and Spider, together with survey work, study how language maps to executable queries under real schemas Zhong2017WikiSQL ; Yu2018Spider ; Ji2022SurveyTextToSQL . Recent systems also apply LLMs to schema matching and metadata tasks, which can reduce the friction of connecting messy enterprise tables to downstream computation Parciak2025LLMMatcher ; mohammed2025entity.
@@
Finance-oriented benchmarks connect these threads. TAT-QA and FinQA require models to retrieve evidence from text and tables and then perform deterministic calculations. Many high-performing approaches use program sketches, tool calls, or verifiable execution traces, which improves auditability and calibration Zhu2021TATQA ; Chen2021FinQA ; mandalawi2025policy. Broader studies of foundation-model use in enterprise settings report promise for back-office tasks, while also warning about hallucinations and schema drift without guardrails and validation Wornow2024AutomatingEnterprise ; althaf2025multi. Work that bridges LLMs and financial analysis reinforces the value of combining language understanding with robust, checkable computation paths LopezLira2025BridgingLMFinance .

### v2

@@ 1 Introduction
Large language models (LLMs) have moved quickly from research demos to everyday tools. They follow instructions, summarize long documents, and support search and analysis at scale. With each generation, they show a stronger grasp of meaning in context, which makes them attractive for business workflows that depend on careful reading and consistent decisions. This momentum has encouraged organizations to pilot LLMs across back-office operations, customer support, and analytics. The same momentum also raises basic questions about reliability. High-stakes applications requires clear standards for transparency, accountability, and risk management, and there is ongoing evidence that fluent outputs can still be incorrect 2; 18; Z. Ji, N. Lee, R. Frieske, T. Yu, D. Su, Y. Xu, E. Ishii, Y. Bang, D. Chen, W. Dai, H. S. Chan, A. Madotto, and P. Fung (2023); S. Qi, L. Gui, Y. He, and Z. Yuan (2024).
@@ 2 Literature Review
Early evaluations showed that strong LLMs still make arithmetic and logic mistakes on word problems. The GSM8K and MATH benchmarks made these issues visible and reproducible Cobbe et al. (2021); Hendrycks et al. (2021). Follow-up studies improved results by asking models to show their steps. Chain-of-Thought (CoT) prompting and Self-Consistency reduce single-path failures and often raise accuracy Wei et al. (2022); Wang et al. (2022). Another line of work reduces mistakes by letting models write and run code, or by calling tools. Program-Aided LM (PAL) and Program-of-Thoughts separate natural language reasoning from exact computation. This improves faithfulness on math and finance-style questions Gao et al. (2022); Chen and others (2022). Domain-adapted models such as Minerva also help on STEM problems, but they still benefit from careful instructions and structure Lewkowycz et al. (2022).
@@
Reasoning strategies continue to evolve. Least-to-Most prompting breaks a problem into smaller parts, which can help generalization to harder cases Zhou et al. (2022). Prompt engineering practices have become more systematic, including patterns that set roles, constraints, and output formats. Toolformer further shows that models can learn when to call external tools such as calculators and search Schick et al. (2023). Recent work also probes arithmetic reliability directly and designs tests that expose subtle numerical errors, which is important when results must be exact Sun and others (2025); Talburt et al. (2026).
@@
Beyond raw reasoning, many practical tasks require models to align text with tables, spreadsheets, or databases. This alignment is critical in payroll, where each field has a specific meaning and must be used in the right order. In table question answering, TaPas learns table representations that support QA without explicit logical forms, and TAPEX pretrains a neural executor for table operations Herzig et al. (2020); Liu et al. (2021). For database-backed settings, large text-to-SQL datasets such as WikiSQL and Spider, together with survey work, study how language maps to executable queries under real schemas Zhong et al. (2017); Yu et al. (2018); Qin et al. (2022). Recent systems also apply LLMs to schema matching and metadata tasks, which can reduce the friction of connecting messy enterprise tables to downstream computation Parciak and others (2025); Mohammed et al. (2025).
@@
Finance-oriented benchmarks connect these threads. TAT-QA and FinQA require models to retrieve evidence from text and tables and then perform deterministic calculations. Many high-performing approaches use program sketches, tool calls, or verifiable execution traces, which improves auditability and calibration Zhu et al. (2021); Chen et al. (2021); Mandalawi et al. (2025). Broader studies of foundation-model use in enterprise settings report promise for back-office tasks, while also warning about hallucinations and schema drift without guardrails and validation Wornow and others (2024); Althaf et al. (2025). Work that bridges LLMs and financial analysis reinforces the value of combining language understanding with robust, checkable computation paths Lopez-Lira and others (2025).